OpenStudy (anonymous):

How do you convert 0.618 km to cm

OpenStudy (calculusxy):

multiply 0.618*100000 or move the decimal five places to the right

OpenStudy (mathstudent55):

1 km = 1000 m 1 m = 100 cm \(0.618~km \times \dfrac{1000~m}{km} \times \dfrac{100~cm}{m} \) The units cancel out leaving cm: \(0.618~\cancel{\color{red}{km}}\times \dfrac{1000~\cancel{\color{blue}{m}}}{\cancel{\color{red}{km}}} \times \dfrac{100~cm}{\cancel{\color{blue}{m}}} \)